Speeding: On a stretch of Interstate-89, car speed is a normally distributed variable with a mean of 68.9 mph and a standard deviation of 4.1 mph.

You are traveling at 75 mph. Approximately what percentage of cars are traveling faster than you? Enter your answer as a percentage with 1 decimal place.
%
Answer #1

Here, μ = 68.9, σ = 4.1 and x = 75. We need to compute P(X >= 75). The corresponding z-value is calculated using Central Limit Theorem

z = (x - μ)/σ
z = (75 - 68.9)/4.1 = 1.49

Therefore,
P(X >= 75) = P(z <= (75 - 68.9)/4.1)
= P(z >= 1.49)
= 1 - 0.9319 = 0.0681
= 6.8%